A goldsmith has 80g of an alloy that is 50% pure gold. How many grams of an alloy that is 80% pure gold must be combined with the 80g of alloy that is 50% pure gold to form an alloy that is 72% pure gold?
Answer by mananth(16946)

------ Percent ----------------quantity
Alloy 1 50 ----------------80
Alloy 2 80 ----------------x
Total 72 ----------------80+x
50*80+80*x=72(80+x)
4000+80x=5760+72x
80x-72x=5760-4000
8x=1760
/8
x=220 gms 80 % Alloy 2